OpenStudy (anonymous):

There are 10 marbles in a bag, and the marbles are either red or blue. Eric will randomly choose two marbles from the bag, without replacing the first one. If the probability of both marbles' being red is 2/15 , how many BLUE marbles are in the bag? A) 3 B) 4 C) 5 D) 6

OpenStudy (anonymous):

number of red marbles in bag = r. b = 10-r Probability of first marble being red: r/10 Second marble being red: (r-1)/9 total probability = 2/15 = (r-1)*r/90 r = 4 b = 6